\begin{thebibliography}{99}

\bibitem{Fourier} de Stainville, Janot (1815). \emph{Mélanges d'Analyse Algébrique et de Géométrie} [A mixture of Algebraic Analysis and Geometry]. Veuve Courcier. pp. 340–341.

\bibitem{Euler1}  Euler, Leonhard (1744). \emph{De fractionibus continuis dissertatio} [A dissertation on continued fractions] (PDF). Commentarii academiae scientiarum Petropolitanae.\textbf{ 9}: 98–137.

\bibitem{Euler2}  Euler, Leonhard (1985). \emph{An essay on continued fractions}. Mathematical Systems Theory., \textbf{18}: 295–398. doi:10.1007/bf01699475. hdl:1811/32133.

\bibitem{Euler3}   Sandifer, C. Edward (2007). \emph{Chapter 32: Who proved e is irrational?}. How Euler did it. Mathematical Association of America. pp. 185–190. ISBN 978-0-88385-563-8. LCCN 2007927658.

\bibitem{Cohn}  Cohn, Henry (2006). \emph{A short proof of the simple continued fraction expansion of e}. American Mathematical Monthly. Mathematical Association of America, \textbf{113 (1)}: 57–62. arXiv:math/0601660. doi:10.2307/27641837. JSTOR 27641837.

\bibitem{MacDivitt} MacDivitt, A. R. G.; Yanagisawa, Yukio (1987), \emph{An elementary proof that e is irrational}, The Mathematical Gazette, London: Mathematical Association, \textbf{71 (457)}: 217, doi:10.2307/3616765, JSTOR 3616765

\bibitem{Penesi}  Penesi, L. L. (1953). \emph{Elementary proof that e is irrational}. American Mathematical Monthly. Mathematical Association of America. \textbf{60 (7)}: 474. doi:10.2307/2308411. JSTOR 2308411.

\bibitem{Apostol}  Apostol, T. (1974).\emph{ Mathematical analysis} (2nd ed., Addison-Wesley series in mathematics). Reading, Mass.: Addison-Wesley.

\bibitem{Liouville1} Liouville, Joseph (1840).  \emph{Sur l'irrationalité du nombre e = 2,718…}. Journal de Mathématiques Pures et Appliquées. 1 (in French). \textbf{5}: 192.

\bibitem{Hurwitz} Hurwitz, Adolf (1933) [1891]. \emph{Über die Kettenbruchentwicklung der Zahl e}. Mathematische Werke (in German). 2. Basel: Birkhäuser. pp. 129–133.

\bibitem{Liouville2} Liouville, Joseph (1840). \emph{Addition à la note sur l'irrationnalité du nombre e}. Journal de Mathématiques Pures et Appliquées. 1 (in French). \textbf{5}: 193–194.

\bibitem{Niven} Niven, Ivan. \emph{Irrational Numbers. 1st ed.}, vol. 11, Mathematical Association of America, 1985. JSTOR, www.jstor.org/stable/10.4169/j.ctt5hh8zn.

\bibitem{Aigner} Aigner, Martin; Ziegler, Günter M. (1998), \emph{Proofs from THE BOOK (4th ed.)}, Berlin, New York: Springer-Verlag, pp. 27–36, doi:10.1007/978-3-642-00856-6, ISBN 978-3-642-00855-9.

\bibitem{Chrystal} G.~Chrystal, \emph{Algebra: An Elementary Text-book for the Higher Classes of Secondary Schools and for Colleges Vol.II} (1889).

\bibitem{Wall} Wall, H. S. \emph{Analytic Theory of Continued Fractions},  The New York: Chelsea, pp. 335-361, particularly `The Continued Fraction of Gauss', 1948.

\bibitem{Borwein}
Borwein, J.; Bailey, D.; and Girgensohn, R. \emph{ Experimentation in Mathematics: Computational Paths to Discovery}. Wellesley, MA: A K Peters, pp. 31-34, particularly `Gauss's Continued Fraction' in §1.8.3 i, 2004.



\end{thebibliography}

\documentclass[10pt,reqno]{amsart}
\usepackage{latexsym}
\usepackage{graphicx}
\usepackage{fancyhdr,amssymb}

\theoremstyle{plain}
\numberwithin{equation}{section}

\theoremstyle{definition}

\begin{document}
\fancyhead{}
\renewcommand{\headrulewidth}{0pt}
\fancyfoot{}
\fancyfoot[LE,RO]{\medskip \thepage}
\fancyfoot[LO]{\medskip The Mathematical Gazette     }
\fancyfoot[RE]{\medskip The Mathematical Gazette}

\setcounter{page}{1}

\title[Another Proof of $e^{x/y}$ being irrational]{Another Proof of $e^{x/y}$ being irrational}
\author{Sourangshu Ghosh}
\address{Department of Civil Engineering\\
                Indian Institute of Technology Kharagpur\\
                Kharagpur,West bengal\\
                India}
\email{sourangshug123@gmail.com   (Corresponding Author)}

\begin{abstract}
    Continued fractions are used to give an alternate proof of $e^{x/y}$ is irrational.
\end{abstract}

\maketitle

\section{Background}
The most well known proof of the irrationality of $e$ was  provided by Joseph Fourier using proof by contradiction \cite{Fourier}. In 1737, Euler gave the first proof of irrationality of $e$ by using the simple continued fraction expansion of $e$ back \cite{Euler1}\cite{Euler2}\cite{Euler3}.

$$e=2+\frac{{1}}{1 + \frac{1}{2 + \frac{1}{1 + \frac{1}{1+\frac{1}{4+\frac{1}{...}}}}}}. \\$$

A simpler proof of this continued fraction was given by Cohn \cite{Cohn}. There are other proofs of $e$ being irrational in the literature such as those by MacDivitt \cite{MacDivitt}, Penesi \cite{Penesi}, Apostol \cite{Apostol}. Higher powers of $e$ were subsequently also proven to be irrational.  The irrationality of $e^{2}$ was proven in \cite{Liouville1} , of  $e^{3}$  in \cite{Hurwitz}, and of $e^{4}$  in \cite{Liouville2}.

A more generalized result where the power is a rational number was proven by Niven in 1985. Ivan Niven \cite{Niven} showed that $e^{x/y}$ is an irrational number using Niven's Polynomials of the form $\frac{x^n(1-x)^n}{n!}$, which can be also be used to prove that $\pi$ is an irrational number. A similar proof was also given by Aigner \cite{Aigner}.

In this document we use continued fraction to give yet another, but simpler, proof of $e^{x/y}$ being irrational, where $x,y$ are integers.

\noindent\rule{0.84in}{0.4pt} \par
\medskip
\indent\indent {\fontsize{8pt}{9pt} \selectfont DOI: 10.35834/YYYY/VVNNPPP \par}
\indent\indent {\fontsize{8pt}{9pt} \selectfont MSC2020: 11J70,11J72 \par}
\indent\indent {\fontsize{8pt}{9pt} \selectfont Key words and phrases: Irrationality, Generalized Continued Fractions,Exponential \par}

\thispagestyle{fancy}

\vfil\eject
\fancyhead{}
\fancyhead[CO]{\hfill Another Proof of $e^{x/y}$ being irrational}
\fancyhead[CE]{S.~Ghosh  \hfill}
\renewcommand{\headrulewidth}{0pt}

\section{Proof}

We will now state the necessary and sufficient condition for the continued fraction proved by Legendre given in Corollary 3, on page 495, in chapter XXXIV on ~~General Continued Fractions" of Chrystal's Algebra Vol.II \cite{Chrystal} to converge into an irrational number.

\noindent \textbf{Theorem:}
\textit{The necessary and sufficient condition that the continued fraction
$$\frac{b_1}{a_1 + \frac{b_2}{a_2 + \frac{b_3}{a_3 + \dots}}}$$}

\textit{ 
\noindent is irrational is that the values $a_{i}, b_{i}$ are all positive integers, and  there is an integer $n$ such that $|a_i| > |b_i|$  for all $i$ greater than $n$.
}

We start with the Continued Fraction Expansion of the hyperbolic tanh function discovered by Gauss \cite{Wall}\cite{Borwein}

$$\tanh z = \frac{z}{1 + \frac{z^2}{3 + \frac{z^2}{5 + \frac{z^2}{...}}}}. \\$$

Then, if we take $z$ to be $\frac{x}{y}$, where $x$ and $y$ are positive integers, we have
$$\tanh z = \frac{\frac{x}{y}}{1 + \frac{(\frac{x}{y})^2}{3 + \frac{(\frac{x}{y})^2}{5 + \frac{(\frac{x}{y})^2}{...}}}}.== \frac{x}{y + \frac{x^2}{3y + \frac{x^2}{5y + \frac{x^2}{...}}}}. \\$$

\noindent \textbf{Corollary:}
\textit{If $x$ and $y$ are integers, then $tan(x/y)$ is irrational
}

Now as $\tanh r=(e^{r}-e^{-r})/(e^{r}+e^{-r})$, it is obvious that if $e^{r}$ is irrational then so is $\tanh r$. Thus $e^{x/y}$ is irrational.

\end{document}